OpenStudy (anonymous):

simplify (order of operations) 6(squared)-45 (divided) 3 (squared) times 2-6

OpenStudy (anonymous):

i got 20 is this right?

OpenStudy (anonymous):

\[6^{2}-\left( \frac{ 45 }{ 3 } \right)^{2}\times 2-6\] This should be the equation.

OpenStudy (anonymous):

Am I right that it is?

OpenStudy (anonymous):

Or is it just \[3^{2} under 45\]